OpenStudy (anonymous):

√5+1/√5-1 im still a bit confused somebody help please

OpenStudy (mertsj):

Multiply numerator and denominator by sqrt5 + 1

OpenStudy (anonymous):

\[(\sqrt{5}+1)(\sqrt{5}+1)/(\sqrt{5}-1)(\sqrt{5}+1)\] Multiplying, you get: \[(5+2\sqrt{5}+1)/(5-1)\] Which gives you \[(6+2\sqrt{5})/(4)\] which reduces to \[(3+\sqrt{5})/2\]

OpenStudy (anonymous):

i get what you did right there but in the end that can be simplified further should i do it or leave it like that?